define chemical properties of water?
1
Expert's answer
2013-03-13T04:10:15-0400
Chemical properties of water
- has a polar nature
- the molecule of water is a dipole
- can form hydrogen bonds (that's why it has relatively high melting and boiling point temperatures)
- it is a good solvent due to its polarity (ionic and polar substances such as acids, alcohols, and salts are relatively soluble in water, and non-polar substances such as fats and oils are not)
- water is amphoteric: it can be an acid or a base in chemical reactions

When water reacts with an acid, it behaves like a base; when it reacts with the base, it behaves like an acid:

HCl (acid) + H2O (base) = H3O+ + Cl−
NH3 (base) + H2O (acid) = NH4+ + OH−

The oxygen atom in water has two lone pairs, that's why water can also be a Lewis base, or electron pair donor, in reactions with Lewis acids:

H+ (Lewis acid) + H2O (Lewis base) = H3O+

Fe3+ (Lewis acid) + H2O (Lewis base) = [Fe(H2O)6]3+

- When a salt of a weak acid or of a weak base is dissolved in water,
water can hydrolyze the salt:

Na2CO3 + H2O = NaOH + NaHCO3

- water can be a ligand in transition metal complexes. Typically it's a monodentate ligand, that forms only one bond with the central atom
- Water contains hydrogen in oxidation state +1 and oxygen in oxidation state −2. It can oxidize substances with reduction potential below the potential of H+/H2 (hydrides, alkali and alkaline earth metals). But such as aluminum, are oxidized by water as well, but their oxides are not soluble, and the reaction stops because of passivation.

2Na + 2H2O = 2NaOH + H2

- Water also can be oxidized, emitting oxygen gas
4AgF2 + 2H2O = 4AgF + 4HF + O2